# In this paper, we investigate the mixed norm estimates for the operator $T$associated with a dilated plane curve $(ut,u\gamma(t))$, defined by

### v2

@@ Abstract
In this paper, we investigate the mixed norm estimates for the operator $T$associated with a dilated plane curve $(ut,u\gamma(t))$, defined by
[…]
@@
where $r\in[1,\infty]$ and $\omega:=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. These results are sharp, except for certain borderline cases. Additionally, we examine the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{u}^{r}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])$ (time-space) estimates for $T$, which are especially almost sharp when $p=2$.
@@ 1 Introduction
A natural question arises concerning the $L_{x}^{p}(\mathbb{R}^{2})$ to $L_{x}^{q}(\mathbb{R}^{2})$ boundedness of $M_{\gamma}$. It is well-established in various contexts that $M_{\gamma}$ is not $L^{p}\rightarrow L^{q}$ bounded unless $p=q$ (see Liu and Yu [34, Remark 1.3]). However, when the supremum is restricted to $u\in[1,2]$, they derived an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}(\mathbb{R}^{2})$ estimate for $M_{\gamma}$, i.e., the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])$ estimate for $T$. This phenomenon, known as $L^{p}$-improving, is discussed in greater detail in [51, 39, 1].
@@
Our first result addresses Hickman’s problem in two-dimensional case, but with some general plane curves including the parabola $\gamma(t):=t^{2}$.
@@
Assume $\gamma\in C^{N}(0,1]$ with $N\in\mathbb{N}$ sufficiently large, $\lim_{t\rightarrow 0^{+}}\gamma(t)=0$, and $\gamma$ is monotonic on (0,1](0,1]( 0 , 1 ]. Additionally, suppose $\gamma$ satisfies the following conditions:
@@
Then, there exists a positive constant $C$ such that, for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
@@
    $\gamma_{2}(t):=a_{d}t^{d}+a_{d+1}t^{d+1}+\cdots+a_{d+m}t^{d+m}$, where $d\geq 2$, $d\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ , $a_{d}\neq 0$, i.e., $\gamma_{2}$ is a polynomial of degree at least $d$ with no linear term and constant term;
@@
    $\gamma_{5}(t)$ is a smooth function on [0,1][0,1][ 0 , 1 ] satisfying $\gamma(0)=\gamma^{\prime}(0)=\cdots=\gamma^{(d-1)}(0)=0$ and $\gamma^{(d)}(0)\neq 0$, where $d\geq 2$ and $d\in\mathbb{N}$. From Iosevich [22], we know that $\gamma_{5}$ is finite type $d$ at $t=0$. $\gamma_{2}$ and $\gamma_{4}$ provide some special cases of $\gamma_{5}$.
@@
[…]
@@
where $\epsilon$ is a positive constant. Let $\dot{H}_{x}^{s}(\mathbb{R}^{n})$ denote the homogeneous Sobolev space. The following space-time estimate for the half-wave propagator is of particular interest in harmonic analysis:
@@
The estimate (1.2) and its local version have been studied by numerous authors. For instance, when $r=\infty$, (1.2) represents a maximal function estimate for the half-wave propagator, which is associated with the pointwise convergence of solutions to their initial data in $L_{x}^{2}(\mathbb{R}^{n})$. Cowling [11] proved that (1.2) holds for all $q=2$ and $s>\frac{1}{2}$. Rogers and Villarroya [38] later extended this result to all $q\in(2,\infty]$. For endpoint estimates and historical remarks, see Cho, Lee, and Li [10]. When the initial data lies in Sobolev spaces $W_{x}^{s,p}(\mathbb{R}^{n})$, and $q=p$, (1.2) holds for $p\geq\frac{2(n+1)}{n-1}$ and $s>(n-1)\left|\frac{1}{2}-\frac{1}{p}\right|$, as a corollary of Bourgain and Demeter [7]. The two-dimensional case of this estimate can be found in Guth, Wang, and Zhang [16], while an endpoint case was established by Heo, Nazarov, and Seeger [17]. When $r=2$ and $q=p>\frac{2(n+2)}{n}$, (1.2) is closely related to Stein’s square function estimates for Bochner-Riesz operators, as discussed in Lee, Rogers, and Seeger [27]. Additionally, analogous results for Fourier integral operators can be found in [43, 35, 36, 12], and for compact manifolds in Beltran, Hickman, and Sogge [3].
@@
Let $\gamma$ be as defined in Theorem 1.1. Then, there exists a positive constant $C$ such that for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
@@ 3. $C:$
From the definition of $\tilde{s}_{A,0}$ with $p^{\prime}<q<3p^{\prime}$, it is evident that Theorem 1.4 is not always sharp, since $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$ is not a linear combination. However, by interpolating between the vertices of the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$, we may obtain some sharp results. It is natural to investigate what occurs in Theorem 1.4 when $p=2$. By interpolating between the points $(\frac{1}{q},\frac{1}{r})=(\frac{1}{6},\frac{1}{6})$ and $(\frac{1}{4},0)$, we obtain that
@@
holds if $(\frac{1}{q},\frac{1}{r})$ lies on the closed line segment $[(\frac{1}{2},1),(\frac{1}{2},0)]$ or within the open pentagon with vertices $(\frac{1}{6},\frac{1}{6})$, $(\frac{1}{2},1)$, $(\frac{1}{2},0)$, and $(\frac{1}{4},0)$. Moreover, this estimate is sharp except the boundary lines. The range of $(\frac{1}{q},\frac{1}{r})$ is illustrated in Figure 2.
@@
We present the detailed result in Theorem 1.4 with $r=2$. Combining Theorem 1.4 with Remark 4.1, if $\left(\frac{1}{p},\frac{1}{q}\right)$ lies on the open line segment ((0,0),(1,1))((0,0),(1,1))( ( 0 , 0 ) , ( 1 , 1 ) ) or within the open (dashed) pentagon with vertices (0,0)(0,0)( 0 , 0 ), (1,1)(1,1)( 1 , 1 ), $\left(\frac{2}{3},\frac{1}{3}\right)$, $\left(\frac{1}{2},\frac{1}{6}\right)$, and $\left(\frac{1}{4},0\right)$, we have
@@
Conversely, (1.6) holds only if $\left(\frac{1}{p},\frac{1}{q}\right)$ belongs to the closed (black) quadrilateral with vertices (0,0)(0,0)( 0 , 0 ), (1,1)(1,1)( 1 , 1 ), $\left(\frac{2}{3},\frac{1}{3}\right)$, and $\left(\frac{1}{3},0\right)$, see Figure 4. We note that if $\gamma(t):=t^{d}$ with $d\geq 3$, the time-space estimate $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{2}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ established in Theorem 1.4 is essentially sharp. However, if $\gamma(t):=t^{d}$ with $d\in(0,1)\cup(1,2)$, the condition $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$ is redundant.
@@
[…]
defined in (2.9), which is essentially a Fourier integral operator with phase function $x\cdot\xi-u|\xi|$. Thus it is basically a half-wave propagator. These estimates for $F_{j,k}$ generalize the results of Mockenhaupt, Seeger, and Sogge [35, Theorem 3.1], and Lee [26, Corollary 1.5] when we restrict our attention to $P_{k}$.
It is evident that the time-space estimate for $T$ differs from the corresponding space-time estimate. Firstly, the latter is translation-invariant when the temporal integral is evaluated before the spatial integral, but not the former. Secondly, the time-space estimate for $T$ not always can be reduced to the corresponding time-space estimate similar to (1.1). Thirdly, we can only obtain a common estimate
@@
for all $r\in[1,\infty]$, not as the corresponding space-time estimate can be improved directly when $r=1$; see (3.3). Our main tools will be interpolation, a fixed time estimate for $F_{j,k}$ based on Seeger, Sogge and Stein [41, Theorem 2.1], Gao, Liu, Miao and Xi [12, Theorem 1.4] and some results stated in Stein [46, Chapters VIII and IX].
@@
holds if $\max\{1+\omega,4\}<p\leq\infty$ and only if $\max\{1+\omega,3\}\leq p\leq\infty$. Thus the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$ in (1.4) is not sharp. Moreover, we can conjecture that the infimum of $p$ in (1.7) is $\max\{1+\omega,3\}$ by the estimate (2)(2)( 2 ) established in Wolff [52]. From Theorem 1.4, we also note that (1.7) is essentially sharp if $\gamma(t):=t^{d}$ with $d\geq 3$.
@@
Finally, we establish some notation conventions. Throughout this paper, the letter $C$ denotes a positive constant, independent of the essential variables, but whose value may vary from line to line. We use $C_{(u,N,\alpha,\ldots)}$ to denote a positive constant depending on the parameters $u,N,\alpha,\ldots$. The notation $a\lesssim b$ (or $a\gtrsim b$) indicates that there exists a finite positive constant $C$ such that $a\leq Cb$ (or $a\geq Cb$). $a\approx b$ means both $a\lesssim b$ and $b\lesssim a$ hold. For $x\in\mathbb{R}^{n}$ and $r\in(0,\infty)$, $B(x,r):=\{y\in\mathbb{R}^{n}:|x-y|<r\}$ and $B^{\complement}(x,r)$ is its complement in $\mathbb{R}^{n}$. Let $\mathbb{S}^{n-1}$ denote the unit sphere on $\mathbb{R}^{n}$. $\mathcal{S}(\mathbb{R}^{n})$ denotes the collection of all Schwartz functions on $\mathbb{R}^{n}$. $\hat{f}$ and $f^{\vee}$ denote the Fourier transform and inverse Fourier transform of $f$, respectively. For $1<q\leq\infty$, $q^{\prime}$ denotes the adjoint number of $q$, i.e., $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. $\mathbb{Z}_{0}^{-}:=\{0,-1,-2,\ldots\}$, :=$\mathbb{N}:=\{1,2,\ldots\}$, $\mathbb{N}_{0}:=\{0,1,2,\ldots\}$, and $\mathbb{R}^{+}:=(0,\infty)$. For any set $E$, $\chi_{E}$ denotes the characteristic function of $E$.
@@ 2 Preliminaries
([32, Lemma 2.2] or [33, Lemma 2.2]) Let $t\in[\frac{1}{2},2]$ and $\Gamma_{j}(t):=\frac{\gamma(2^{j}t)}{\gamma(2^{j})}$ with $j\in\mathbb{Z}_{0}^{-}$. The following inequalities hold uniformly in $j$:
@@
([42, Lemma 2.4.2]) Suppose that $F$ is $C^{1}(\mathbb{R})$. Then, if $q>1$,
@@
([34, Lemma 2.3]) Recall that $\omega=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. Then, for all $(\frac{1}{p},\frac{1}{q})$ satisfying $\frac{1}{q}\leq\frac{1}{p}$ and $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$, we have
@@
where, without loss of generality, we assume $f\geq 0$. Let $\psi$ be a smooth and positive function supported on the interval $\{t\in\mathbb{R}:\frac{1}{2}\leq t\leq 2\}$ such that $\sum_{j\in\mathbb{Z}}\psi_{j}(t)=1$ for all $t>0$, where $\psi_{j}(t):=\psi(2^{-j}t)$. As in [34, 31], it suffices to show that
@@
where the multipliers of $\widetilde{T}^{0}_{j}$ and $\widetilde{T}_{j,k}$ are $H_{j}(u,\xi)\Psi^{0}(u|\xi|)$ and $H_{j}(u,\xi)\psi_{k}(u|\xi|)$, respectively. For $\widetilde{T}^{0}_{j}$, as in [34, 31], by Lemmas 2.1 and 2.2, Hölder’s inequality, and Young’s inequality, it follows that
@@
respectively. Here, $\rho\in C^{\infty}_{c}(\mathbb{R}^{+})$ is a smooth cutoff function satisfying $\rho=1$ on the interval $[\frac{C^{(1)}_{1}}{10e^{C^{(1)}_{2}}},10e^{C^{(1)}_{2}}C^{(1)}_{2}].$ Then, as in (2.2), by Lemmas 2.1 and 2.2, we obtain the decay estimate
@@
which is a Fourier integral operator of order $-\frac{1}{2}$ and satisfies the special assumption that

### v3

# In this paper, we investigate the mixed norm estimates for the operator TTassociated with a dilated plane curve $(ut,u\gamma(t))$, defined by

@@ Abstract
In this paper, we investigate the mixed norm estimates for the operator TTassociated with a dilated plane curve $(ut,u\gamma(t))$, defined by
$Tf(x,u):=\int_{0}^{1}f(x_{1}-ut,x_{2}-u\gamma(t))\,dt,$
@@
where $r\in[1,\infty]$ and $\omega:=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. These results are sharp, except for certain borderline cases. Additionally, we examine the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{u}^{r}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])$ (time-space) estimates for $T$, which are especially almost sharp when $p=2$ or $p\in[1,\frac{3}{2}]\cup[4,\infty]$.
@@ 1 Introduction
A natural question arises concerning the $L_{x}^{p}(\mathbb{R}^{2})$ to $L_{x}^{q}(\mathbb{R}^{2})$ boundedness of $M_{\gamma}$. It is well-established in various contexts that $M_{\gamma}$ is not $L^{p}\rightarrow L^{q}$ bounded unless $p=q$ (see Liu and Yu [35, Remark 1.3]). However, when the supremum is restricted to $u\in[1,2]$, they derived an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}(\mathbb{R}^{2})$ estimate for $M_{\gamma}$, i.e., the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])$ estimate for $T$. This phenomenon, known as $L^{p}$-improving, is discussed in greater detail in [52, 40, 1, 32].
@@
The main purpose of this paper is to consider the mixed norm estimates for $T$ in two-dimensional case, but with some general plane curves including the parabola $\gamma(t):=t^{2}$. Our first result is the following.
@@
Assume $\gamma\in C^{N}(0,1]$ with $N\in\mathbb{N}$ sufficiently large, $\lim_{t\rightarrow 0^{+}}\gamma(t)=0$, and $\gamma$ is monotonic on $(0,1]$. Additionally, suppose $\gamma$ satisfies the following conditions:
@@
Then, there exists a positive constant $C$ depending only on $\gamma$, $p$, $q$ and $r$, such that for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
@@
    $\gamma_{2}(t):=a_{d}t^{d}+a_{d+1}t^{d+1}+\cdots+a_{d+m}t^{d+m}$, where $\geq 2, dd\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ , $a_{d}\neq 0$, i.e., $\gamma_{2}$ is a polynomial of degree at least $d$ with no linear term and constant term;
@@
    $\gamma_{5}(t)$ is a smooth function on $[0,1]$ satisfying $\gamma(0)=\gamma^{\prime}(0)=\cdots=\gamma^{(d-1)}(0)=0$ and $\gamma^{(d)}(0)\neq 0$, where $d\geq 2$ and $d\in\mathbb{N}$. From Iosevich [22], we know that $\gamma_{5}$ is finite type $d$ at $t=0$. $\gamma_{2}$ and $\gamma_{4}$ provide some special cases of $\gamma_{5}$.
@@
The range of $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$ is illustrated in Figure 1. This range corresponds to the open convex hull of the points $\{\textrm{O, A, B, C, D}\}$ but with the open line segment $(\textrm{O, A})$. From Figure 1, it is evident that the range of $(\frac{1}{p},\frac{1}{q})$ diminishes as $r$ increases. Specifically, when $r=1$, the range encompasses the entire lower triangle except for the boundary lines $\frac{1}{p}=1$ and $\frac{1}{q}=0$. Conversely, when $r=\infty$, the range coincides with the result obtained by Liu and Yu [35]. Additionally, the range of $(\frac{1}{p},\frac{1}{q})$ is observed to shrink as $\omega$ increases. Notably, the condition $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$ does not impose any restrictions on the range of $(\frac{1}{p},\frac{1}{q})$ when $\omega\leq\frac{4(r-1)}{r+4}$.
@@
where $\epsilon$ is a positive constant. It is well known that the solution $g$ of the Cauchy problem for the wave equation in $\mathbb{R}^{n}\times\mathbb{R}$
$\begin{cases}(\partial^{2}_{u}-\Delta)g(x,u)=0,~~(x,u)\in\mathbb{R}^{n}\times\mathbb{R};\\ g(x,0)=f(x),~~\partial_{u}g(x,0)=0,\end{cases}$
can be written in terms of the half-wave propagator $e^{iu\sqrt{-\Delta}}f$ as
$g(x,u)=\frac{1}{2}\left(e^{iu\sqrt{-\Delta}}f(x)+e^{-iu\sqrt{-\Delta}}f(x)\right).$
Let $\dot{H}_{x}^{s}(\mathbb{R}^{n})$ denote the homogeneous Sobolev space. The following space-time estimate for the half-wave propagator is of particular interest in harmonic analysis:
@@
The estimate (1.2) and its local version have been studied by numerous authors. For instance, when $r=\infty$, (1.2) represents a maximal function estimate for the half-wave propagator, which is associated with the pointwise convergence of solutions to their initial data in $L_{x}^{2}(\mathbb{R}^{n})$. Cowling [11] proved that (1.2) holds for all $q=2$ and $s>\frac{1}{2}$. Rogers and Villarroya [39] later extended this result to all $q\in(2,\infty]$. For endpoint estimates and historical remarks, see Cho, Lee, and Li [10]. When the initial data lies in Sobolev spaces $W_{x}^{s,p}(\mathbb{R}^{n})$, and $q=p$, (1.2) holds for $p\geq\frac{2(n+1)}{n-1}$ and $s>(n-1)\left|\frac{1}{2}-\frac{1}{p}\right|$, as a corollary of Bourgain and Demeter [7]. The two-dimensional case of this estimate can be found in Guth, Wang, and Zhang [16], while an endpoint result in dimensions $n\geq 4$ was established by Heo, Nazarov, and Seeger [17]. When $r=2$ and $q=p>\frac{2(n+2)}{n}$, (1.2) is closely related to Stein’s square function estimates for Bochner-Riesz operators, as discussed in Lee, Rogers, and Seeger [27]. Additionally, analogous results for Fourier integral operators can be found in [44, 36, 37, 12], and for compact manifolds in Beltran, Hickman, and Sogge [3].
@@
Let $\gamma$ be as defined in Theorem 1.1. Then, there exists a positive constant $C$ depending only on $\gamma$, $p$, $q$ and $r$, such that for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
@@ 3. $C:$
From the definition of $\tilde{s}_{A,0}$ with $p^{\prime}<q<3p^{\prime}$, it is evident that Theorem 1.4 is not always sharp, since the indices $\frac{1}{p}$, $\frac{1}{q}$ and $\frac{1}{r}$ in $\tilde{s}_{A,0}$ is not a linear combination. However, by interpolating between the vertices of the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$, we may obtain some sharp results. It is natural to investigate what occurs in Theorem 1.4 when $p=2$. By interpolating between the points $(\frac{1}{q},\frac{1}{r})=(\frac{1}{6},\frac{1}{6})$ and $(\frac{1}{4},0)$, we obtain that
@@
holds if $(\frac{1}{q},\frac{1}{r})$ lies on the closed line segment $(\frac{1}{2},1),(\frac{1}{2},0)]$ or within the open pentagon with vertices $(\frac{1}{6},\frac{1}{6})$, $(\frac{1}{2},1)$, $(\frac{1}{2},0)$, and $(\frac{1}{4},0)$. Moreover, this estimate is sharp except the boundary lines. The range of $(\frac{1}{q},\frac{1}{r})$ is illustrated in Figure 2.
@@
We present the detailed result in Theorem 1.4 with $r=2$. Combining Theorem 1.4 with Remark 4.1, if $\left(\frac{1}{p},\frac{1}{q}\right)$ lies on the open line segment $((0,0),(1,1))$ or within the open (dashed) pentagon with vertices $(0,0)$, $(1,1)$, $\left(\frac{2}{3},\frac{1}{3}\right)$, $\left(\frac{1}{2},\frac{1}{6}\right)$, and $\left(\frac{1}{4},0\right)$, we have
@@
Conversely, (1.6) holds only if $\left(\frac{1}{p},\frac{1}{q}\right)$ belongs to the closed (black) quadrilateral with vertices $(0,0)$, $(1,1)$, $\left(\frac{2}{3},\frac{1}{3}\right)$, and $\left(\frac{1}{3},0\right)$, see Figure 4. We note that if $\gamma(t):=t^{d}$ with $d\geq 3$, the time-space estimate $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{2}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ established in Theorem 1.4 is essentially sharp. However, if $\gamma(t):=t^{d}$ with $d\in(0,1)\cup(1,2)$, the condition $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$ is redundant.
@@
$F_{j,k}f(x,u):=\Omega(x,u)\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}e^{i\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,\xi)\hat{f}(\xi)\,d\xi,$
defined in (2.9), which is essentially a Fourier integral operator with phase function $x\cdot\xi-u|\xi|$. Thus it is basically a half-wave propagator. These estimates for $F_{j,k}$ generalize the results of Mockenhaupt, Seeger, and Sogge [36, Theorem 3.1], and Lee [26, Corollary 1.5] when we restrict our attention to $P_{k}$.
It is evident that the time-space estimate for $T$ differs from the corresponding space-time estimate. Firstly, the latter is translation-invariant when the temporal integral is evaluated before the spatial integral, but not the former. Secondly, as stated in Subsection 4.1, the time-space estimate for $T$ not always can be reduced to the corresponding time-space estimate similar to (1.1). Thirdly, we can only obtain a common estimate
@@
for all $r\in[1,\infty]$, not as the corresponding space-time estimate can be improved directly when $r=1$; see (3.3). Our main tools will be interpolation, a fixed time estimate for $F_{j,k}$ based on Seeger, Sogge and Stein [42, Theorem 2.1], Gao, Liu, Miao and Xi [12, Theorem 1.4] and some results stated in Stein [47, Chapters VIII and IX].
@@
holds if $\max\{1+\omega,4\}<p\leq\infty$ and only if $\max\{1+\omega,3\}\leq p\leq\infty$. Thus the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$ in (1.4) is not sharp. Moreover, we can conjecture that the infimum of $p$ in (1.7) is $\max\{1+\omega,3\}$ by the estimate $(2)$ established in Wolff [53]. From Theorem 1.4, we also note that (1.7) is essentially sharp if $\gamma(t):=t^{d}$ with $d\geq 3$.
@@
Finally, we establish some notation conventions. Throughout this paper, the letter $C$ denotes a positive constant, independent of the essential variables, but whose value may vary from line to line. We use $C_{(u,N,\alpha,\ldots)}$ to denote a positive constant depending on the parameters $u,N,\alpha,\ldots$. The notation $a\lesssim b$ (or $a\gtrsim b$) indicates that there exists a finite positive constant $C$ such that $a\leq Cb$ (or $a\geq Cb$). $a\approx b$ means both $a\lesssim b$ and $b\lesssim a$ hold. For $x\in\mathbb{R}^{n}$ and $r\in(0,\infty)$, $B(x,r):=\{y\in\mathbb{R}^{n}:|x-y|<r\}$ and $B^{\complement}(x,r)$ is its complement in $\mathbb{R}^{n}$. Let $\mathbb{S}^{n-1}$ denote the unit sphere on $\mathbb{R}^{n}$. $\mathcal{S}(\mathbb{R}^{n})$ denotes the collection of all Schwartz functions on $\mathbb{R}^{n}$. $\hat{f}$ and $f^{\vee}$ denote the Fourier transform and inverse Fourier transform of $f$, respectively. For $1<q\leq\infty$, $q^{\prime}$ denotes the adjoint number of $q$, i.e., $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. $\mathbb{Z}_{0}^{-}:=\{0,-1,-2,\ldots\}$, $\mathbb{N}:=\{1,2,\ldots\}$, $\mathbb{N}_{0}:=\{0,1,2,\ldots\}$, and $\mathbb{R}^{+}:=(0,\infty)$. For any set $E$, $\chi_{E}$ denotes the characteristic function of $E$.
@@ 2 Preliminaries
([33, Lemma 2.2] or [34, Lemma 2.2]) Let $\gamma$ be defined as in Theorem 1.1 and let $\Gamma_{j}(t):=\frac{\gamma(2^{j}t)}{\gamma(2^{j})}$ with $j\in\mathbb{Z}_{0}^{-}$ and $t\in[\frac{1}{2},2]$. The following inequalities hold uniformly in $j$:
@@
([43, Lemma 2.4.2]) Suppose that $F$ is $C^{1}(\mathbb{R})$. Then, if $q>1$,
@@
([35, Lemma 2.3]) Recall that $\omega=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. Then, for all $(\frac{1}{p},\frac{1}{q})$ satisfying $\frac{1}{q}\leq\frac{1}{p}$ and $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$, we have
@@
where, without loss of generality, we assume $f\geq 0$. Let $\psi$ be a smooth and positive function supported on the interval $\{t\in\mathbb{R}:\frac{1}{2}\leq t\leq 2\}$ such that $\sum_{j\in\mathbb{Z}}\psi_{j}(t)=1$ for all $t>0$, where $\psi_{j}(t):=\psi(2^{-j}t)$. For any $j\in\mathbb{Z}$, we define $\delta_{j}f(x_{1},x_{2}):=f(2^{j}x_{1},\gamma(2^{j})x_{2})$. Note that $|2^{j}\gamma(2^{j})|^{\frac{1}{q}}\|\delta_{j}f\|_{L^{q}(\mathbb{R}^{2})}=\|f\|_{L^{q}(\mathbb{R}^{2})}$, it suffices to show that
@@
where the multipliers of $\widetilde{T}^{0}_{j}$ and $\widetilde{T}_{j,k}$ are $H_{j}(u,\xi)\Psi^{0}(u|\xi|)$ and $H_{j}(u,\xi)\psi_{k}(u|\xi|)$, respectively. For $\widetilde{T}^{0}_{j}$, by Lemma 2.1, we can obtain that $|\partial_{u}^{\alpha}\partial^{\beta}_{\xi}(H_{j}(u,\xi)\Psi^{0}(u|\xi|))|\lesssim(1+|\xi|)^{-3}$ for all $(\alpha,\beta)\in\mathbb{N}_{0}\times\mathbb{N}^{2}_{0}$ with $|\alpha|\leq 1$ and $|\beta|\leq 3$, where the implicit constant is independent of $u\in[1,2]$ and $j\in\mathbb{Z}_{0}^{-}$. Furthermore, by Lemma 2.2, Hölder’s inequality, and Young’s inequality, it follows that
@@
respectively. Here, $\rho\in C^{\infty}_{c}(\mathbb{R}^{+})$ is a smooth cutoff function satisfying $\rho=1$ on the interval $\frac{C^{(1)}_{1}}{10e^{C^{(1)}_{2}}},10e^{C^{(1)}_{2}}C^{(1)}_{2}$]. Then, as in (2.2), by Lemmas 2.1 and 2.2, we obtain the decay estimate
@@
which is a Fourier integral operator of order $\frac{1}{2}$ and satisfies the special assumption that